OpenStudy (destinyyyy):

I need help on 4 Exponential and Logarithmic Equations..

OpenStudy (destinyyyy):

Solve the exponential equation. Express the solution in terms of natural logarithms. Then, use a calculator to obtain a decimal approximation for the solution. e^5-3x =1210

OpenStudy (destinyyyy):

I got x= 5-Ln1210 over 3 = 2.63

OpenStudy (destinyyyy):

Its telling me 2.63 is wrong.

OpenStudy (destinyyyy):

Yes.

OpenStudy (destinyyyy):

Um the correct answer for the first part is x= 5-Ln1210 over 3

OpenStudy (destinyyyy):

I need help with the second part. The decimal approximation.

OpenStudy (anonymous):

You have: \[e^{5-3x}=1210\] Now take the reciporicals on both sides: \[e^{3x-5}=\frac{ 1 }{ 1210 }\] Now you want to eliminate the exponential from the left hand side. You take the natural logarithm on both sides: \[3x-5=-log(1210)\] Solve x: \[3x=5-log(1210)\] \[x=\frac{5-log(1210)}{3}\]

OpenStudy (destinyyyy):

Your answer is incorrect. I already said what the correct answer is to the first part. I was asking for help on the second part.